# Relative Overfitting and Accept-Reject Framework

### v1

@@ 4.2 AR Law
The AR law constitutes the rule within the AR framework that defines the acceptance and rejection criteria for judgments of various SLMs when operating with a fixed LLM. This chapter will illustrate this based on the ideas, discussions, and experimental results of our approach, using the simple implementation under the AR framework described in Section 4.1 as an example.
First, we introduce the concept of the threshold point of AR. Let the parameter ratio between the LLM and the SLM be $k=\frac{|\theta_{\mathrm{LLM}}|}{|\theta_{\mathrm{SLM}}|}$, where $|\theta_{\mathrm{LLM}}|$ denotes the number of parameters in the LLM, and $|\theta_{\mathrm{SLM}}|$ denotes the number of parameters in the SLM. Since the performance of the same model can vary across different datasets, for convenience, this paper uses parameter count as a proxy for the concept of performance. It is important to note that although our discussion uses the parameter ratio as a metric, it represents the performance ratio and the underlying concept of relative overfitting. Therefore, if the AR weight allocation ratio that achieves the optimal performance improvement is $\alpha$, then we have:
@@
Second, we introduce the concept of the improvement domain, denoted as $[R,\infty]$, where $R<\alpha$. For any weight allocation ratio $\beta$, when $\beta\in[R,\infty]$, the model can achieve a performance improvement. Furthermore, when $\beta<\alpha$, the magnitude of performance improvement increases, and when $\beta>\alpha$, the magnitude of performance improvement decreases. When $\beta=\infty$, it approximates a single model. We also have:
@@
When $k$ is extremely large, meaning the performance (parameter) gap between the models is vast, $R$ approaches infinity ($R\to\infty$). In this scenario, we consider the performance improvement to be near zero. Of course, this situation rarely occurs. To verify this extreme case, we designed an experiment using a model with extremely poor performance and very high perplexity to construct the AR framework. In this case, the performance improvement was almost negligible.
@@
The points described above constitute the basic AR law defined on the basis of the theoretical ideas presented earlier. Next, we will discuss the selection of the threshold point of AR.
@@ 4.3 The Threshold Point of AR
Based on the ideas discussed in the previous section, we examine the optimal allocation ratio threshold point $\alpha$ based on experimental results. We consider two main influencing factors in the entire system:
  • Marginal effect of the performance gap between LLM and SLM: In most cases, we tend to leverage details, meaning that as the performance gap widens, the LLM should be assigned a greater weight.
  • Relative overfitting: Relative overfitting tends to shift the leverage weight ratio slightly towards the SLM.
We decompose the optimal ratio $\alpha$ into two conflicting components. The first is an ascending component. If the parameter ratio $k$ between the LLM and SLM is large, we should use a larger $\alpha$ to better utilize the LLM’s capability for detail, essentially, its conventionally understood performance. According to Kaplan 1 , LLM performance follows a power law concerning parameter count, suggesting($p$ is the power-law exponent):
@@
The second component characterizes the magnitude at which the AR framework operates effectively. In reality, $k$ only describes the gap between the LLM and SLM but fails to capture the starting point of this gap. We introduce the parameter $\theta_{LLM}$ (the LLM’s parameter count) to represent the starting point where this influence takes effect. For example, the ratio between parameters 14B, 7B, 14M, and 28M is also 2, but the magnitude of the differences is entirely different. The relative overfitting cases arising from different magnitudes will also vary. Therefore, as noted in Section 4.2, we are discussing how to influence it based on the AR framework and AR law under a fixed LLM. The average optimal weight $\alpha$ of each model across various datasets in all experiments from Section 5.1 is presented in Appendix B.2.
@@ 4.4 Improvement Magnitude
In this subsection, we discuss the magnitude of performance improvement contributed by the lower-performing model to the higher-performing model within the AR framework. Differing from previous discussions, the magnitude of performance improvement, while dependent on the weights, is influenced by more complex factors in practice than the relatively strict proportional and inverse relationships explored in Section 4.2. Based on experimental results, we identify two empirical regularities. First, as the weight ratio (associated with the LLM) increases, the proportion and influence of the SLM diminish. However, this inverse relationship is not absolute; exceptions were observed in experiments. This is because the change in proportion is an integral part of the AR law itself—reducing the SLM’s proportion serves the mechanism’s requirements and is not merely about decreasing the SLM’s influence. we see that the actual optimal weight ratio $\alpha$ increases only with $k$ (the parameter ratio).
Secondly, as stated in Section 4.3. We observe that while the actual optimal weight ratio $\alpha$ increases with our experiments indicate it does not demonstrably increase with the scale of the main model parameters $\theta_{LLM}$. This provides evidence for the stability of the AR framework: namely, as long as an appropriate k is maintained, the increase in the model’s performance remains stable with the growth of $\theta_{LLM}$, and in some scenarios, there is even a rising trend (as seen in Section 5.1). This means that, according to the AR law, we need not worry about the significant reduction in performance due to excessively large main model parameters. This finding supports our belief that this approach can positively impact the bottleneck observed in scaling laws. And under the scaling law, as the parameters of LLM increase, the rate of improvement decreases rapidly.
@@ 5 Experiments
To validate the method proposed in this paper, in the field of NLP, we designed experiments considering existing mainstream scaling-law-based LMs as well as custom-built LMs based on prevalent architectures. The aim is to verify the universality, effectiveness, stability, and flexibility of our approach. First, we tested our method on several existing scaling-law-based model series (e.g., GPT-2 33 ). Second, we conducted ablation studies using a model built upon the Transformer 31 architecture, deliberately avoiding any task-specific design modifications, and comparing against existing models. Finally, we retrained models of varying scales based on the mainstream architectures LSTM 29 ; 30 , Transformer, and GRU 32 for further experimentation. In summary, this chapter demonstrates that the proposed method is universally applicable across a variety of models, including mainstream models, custom-built models, and combinations thereof. Furthermore, we also validated the AR framework, based on the principle of relative overfitting, in the domains of computer vision (CV) and AI for science.
@@ 5.1 Mainstream Language Model Testing
This paper conducted experiments using the proposed AR framework on several pre-trained models. Tests were performed using GPT-2 33 on fundamental language modeling benchmarks including WikiText2, WikiText103 34 , text8 14 , Pile 39 ; 40 (using a randomly sampled subset of the Pile test set), and 1BW 38 , as well as on the long-context benchmark LAMBADA 35 and QA benchmarks CBT 36 and subject examination benchmark ARC 37 . We also evaluated the performance of models with different architectures—Pythia 41 (multi-scale scaling law), RWKV 42 (combining RNN and Transformer), Mamba 43 (SSM series), and Qwen 44 ; 45 (commercial LLM)—using the AR framework for basic language modeling tasks. To compare the behavior under the AR framework between general language modeling and specific tasks, we used the LAMBADA dataset for both language modeling evaluation (perplexity) and its original long-context dependency evaluation.
The results in Table 1 demonstrate that through our AR framework, all models achieve universal performance improvements. In many scenarios, we can achieve superior performance with significantly fewer parameters and less computation time. For instance, on LAMBADA (PPL), Pile, and other benchmarks, our method consistently outperforms the strategy of simply increasing the LLM’s parameter count, achieving superiority in terms of parameters, computation time, and performance. This advantage becomes more pronounced as the model parameters increase. For example, based on the improvement rate observed when scaling GPT-2 from 762M to 1.5B, and roughly estimating the improvement from 1.5B to 3B using the power-law model proposed in scaling laws, the performance of the 1.5B+762M combination almost universally surpasses the projected 3B model across all benchmarks. Notably, our method’s advantages become more significant with larger models, while the linear complexity related to vocabulary size becomes negligible compared to increasing the parameter count. Furthermore, as the main model’s parameters grow, the magnitude of performance improvement increases rather than diminishes, leading us to be optimistic about its performance on even larger models and hopeful that it can contribute to alleviating the scaling law bottleneck.
Table 1: The table shows the test results of the classic model GPT2 based on the AR framework with four parameter sizes from 117M to 1.5B. LTD is the long context benchmark under the LAMBADA dataset. The results of RWKV, Mamba, Pythia, and Qwen can be found in the appendix.
@@ 5.2 Ablation Experiments with Custom-Built Models Based on Mainstream Architectures
To verify the universality and flexibility of our proposed method, in this section, we utilized the vocabularies corresponding to Pythia, RWKV, Mamba, and GPT-2. Without incorporating any specific design elements, we employed a simple Transformer architecture to train a reference model with only approximately 180M parameters. In practice, the standalone performance of this model is not particularly strong.
To further validate the universality, stability, and flexibility of our method, we designed a control experiment. On the WikiText103 and LAMBADA datasets, we performed predictions using only the model trained on WikiText2, without using the corresponding training sets for these target datasets. This was done to confirm that the performance improvement achieved by the AR framework is not merely due to having "seen the data". Additionally, to verify the effectiveness of our method, we trained models on the 1BW and text8 datasets using their respective training sets, aiming to achieve substantial performance gains.
The results in Table 2 indicate that although the models we trained on WikiText2 and WikiText103 did not perform well individually, their improvement magnitudes exceeded those of inherently superior but highly homogeneous models. The results on the LAMBADA dataset further demonstrate that even extremely poor-performing models can still contribute improvements through the AR framework. The outcomes on the text8 and 1BW datasets illustrate that our method is not only universal but also highly effective (even though we did not specifically search for the optimal fulcrum $\alpha$). The AR framework based on non-homogeneous models has nearly achieved comprehensive advantages in computation, parameters, and performance in certain scenarios compared to LLMs that seek improvement by increasing the number of parameters. Even without additional design, enhancements will occur as long as an appropriate threshold point is identified based on relative overfitting using the AR framework.
It is important to note that this is not fine-tuning 46 ; 47 , but merely a training-free bias correction. Compared to fine-tuning, our method offers greater flexibility, requires significantly fewer resources, and does not affect the base model (domain-specific fine-tuning may lead to performance degradation in other domains, as seen with DeepSeekR1 48 ).
@@
Figure 2: Test results of AR framework based on simple LSTM, Transformer and GRU architectures under different embedding dimensions of SLM and LLM..
@@ 5.3 Self-built model testing based on mainstream architectures
This paper conducted tests based on the current mainstream model architectures LSTM 29 ; 30 , Transformer 31 , and GRU 32 , utilizing the same vocabulary but varying embedding dimensions. In practice, we employed very simple architectures 49 ; 50 ; 51 ; 52 without any specific model design modifications. Even under these conditions, our method continues to demonstrate universal and stable effectiveness (as illustrated in Figure 2).
@@ 5.4 Experiments in the rest of the machine learning field
Based on the principles of relative overfitting and the AR framework, we employed a similar methodology to conduct experiments with our approach across a broader range of machine learning domains, including CV and AI for Science, thereby validating the general effectiveness of our proposed framework (further details are available in Appendix C).
@@ 6 Discussion and Future work
  • Model Selection Trade-offs: For homogeneous scaling models, while performance showed stability (no significant fluctuations) as the main model grew, the improvement magnitude diminished as the parameter ratio expanded. Larger ratios necessitated more extreme leverage, reducing SLM effectiveness and potentially limiting resource savings. Thus, in specific domains, non-homogeneous models might be a better choice.
  • Better Improvement Magnitude: Although this paper proposes the AR law based on relative overfitting and demonstrates its universal, stable, and effective improvements along with good flexibility under a suitable threshold point, significant room for progress remains beyond the decent performance improvements already achieved. To reveal fundamental principles, our method attained these results despite lacking any specific model design, highlighting the potential yet to be unlocked. Therefore, with a deeper understanding of the entire system and relative overfitting, key research questions worth exploring include different combination methods based on the AR law, the selection and innovation of low-parameter reference model architectures serving as backbones, and how to better maintain the information purity of the SLM. We anticipate that addressing these questions will lead to the emergence of greater improvement magnitudes.
@@
  • Broader Applicability: Although this paper primarily focuses on the NLP domain and has also conducted some experiments in CV and AI for Science, due to limited time and resources, a broader discussion across more fields was not feasible. The applicability of the relative overfitting phenomenon and the proposed AR framework in a wider range of domains warrants further exploration.

### v4

@@ 4.2 AR Law
The AR law constitutes the rule within the AR framework that defines the acceptance and rejection criteria for the judgments of various SLMs when operating with a fixed LLM. This chapter will illustrate this based on the ideas, discussions and experimental results of our approach, using the simple implementation of the AR framework described in Section 4.1 as an example.
First, we introduce the concept of the threshold point of AR. Let the parameter ratio between the LLM and the SLM be $k=\frac{|\theta_{\mathrm{LLM}}|}{|\theta_{\mathrm{SLM}}|}$, where $|\theta_{\mathrm{LLM}}|$ denotes the number of parameters in the LLM, and $|\theta_{\mathrm{SLM}}|$ denotes the number of parameters in the SLM. Since the performance of the same model can vary across different datasets, we use parameters as proxy indicators for the inherent attributes of the model. Although our discussion uses the parameter ratio as a metric, it is important to note that it represents the performance ratio and the underlying concept of relative overfitting. We define the weight ratio as $\beta=w/(1-w)$. Therefore, if the AR weight allocation ratio that achieves the optimal performance improvement is $\alpha$ ($\alpha$ is designed to maximize the performance improvement of the ensemble model $\beta$), then we have:
@@
Second, we introduce the concept of the improvement domain, denoted as $[D,\infty]$, where $D<\alpha$. For any weight allocation ratio $\beta$, the model can improve performance when $\beta\in[D,\infty]$. Furthermore, when $\beta<\alpha$, the magnitude of performance improvement increases, and when $\beta>\alpha$, the magnitude of performance improvement decreases. When $\beta=\infty$, it approximates a single model. We also have:
@@
When $k$ is extremely large, meaning the performance (parameter) gap between the models is vast, $R$ approaches infinity ($R\to\infty$). In this scenario, we consider the performance improvement to be near zero. Of course, this situation rarely occurs. To verify this extreme case, we designed an experiment using a model with abysmal performance and very high perplexity to construct the AR framework. In this case, the performance improvement was almost negligible.
@@
The points described above constitute the basic AR law, which is defined on the basis of the theoretical ideas presented earlier. Next, we will discuss the selection of the AR threshold point.
@@ 4.3 The Threshold Point of AR
Based on the ideas discussed in the previous section, we examine the threshold point of the optimal allocation ratio $\alpha$ based on the experimental results.
We decompose the optimal ratio $\alpha$ into two conflicting components. The first is an ascending component. If the parameter ratio $k$ between the LLM and the SLM is large, we should use a larger $\alpha$ to better utilize the LLM’s capability for detail, essentially its conventionally understood performance. According to Kaplan 1 , LLM performance follows a power law concerning parameter count, suggesting($p$ is the power-law exponent):
@@
The second component characterizes the magnitude at which the AR framework operates effectively. In reality, $k$ only describes the gap between LLM and SLM but does not capture the starting point of this gap. We introduce the parameter $\theta_{LLM}$ (the LLM’s parameter count) to represent the starting point where this influence takes effect. For example, the ratio between parameters 14B, 7B, 14M, and 28M is also 2, but the magnitude of the differences is entirely different. Relative overfitting cases arising from different magnitudes will also vary. Therefore, as noted in Section 4.2, we discuss how to influence it based on the AR framework and law under a fixed LLM. The average optimal weight $\alpha$ of each model in various datasets in all experiments from Section 5.1 is presented in Appendix B.2.
@@ 4.4 Improvement Magnitude
In this subsection, we discuss the magnitude of performance improvement contributed by the lower performing model to the higher performing model within the AR framework. Unlike previous discussions, although dependent on weights, the magnitude of performance improvement is influenced by more complex factors in practice than the relatively strict proportional and inverse relationships explored in Section 4.2. Based on the experimental results, we identify two empirical regularities. First, as the weight ratio (associated with the LLM) increases, the proportion and influence of the SLM decrease. However, this inverse relationship is not absolute; exceptions were observed in experiments. This is because the change in proportion is an integral part of the AR law itself: reducing the SLM’s proportion serves the mechanism’s requirements and is not merely about decreasing the SLM’s influence. we see that the actual optimal weight ratio $\alpha$ increases only with $k$ (the parameter ratio). Secondly, as stated in Section 4.3. We observe that while the actual optimal weight ratio $\alpha$ increases with our experiments, it does not demonstrably increase with the scale of the main model parameters $\theta_{LLM}$. This provides evidence for the stability of the AR framework: that is, as long as an appropriate k is maintained, the increase in the model’s performance remains stable with the growth of $\theta_{LLM}$, and in some scenarios, there is even a rising trend (as seen in Section 5.1). This means that, according to the AR law, we need not worry about a significant reduction in performance due to extensive main model parameters. This finding supports our belief that this approach can positively impact the bottleneck observed in scaling laws. Moreover, under the scaling law, as the LLM parameters increase, the rate of improvement decreases rapidly.
4.5 Proof of part of the AR law
Chapters 4.2, 4.3, and 4.4 present empirical conclusions derived from our tens of thousands of experimental results. In fact, we provide a rigorous proof concerning the AR law based on the ACC metric, under specific assumptions, as seen in Appendix A. This proof not only demonstrates a part of the correctness of AR law, but its proof process also more comprehensively illustrates the ideas contained within the AR framework.
In particular, we can naturally extend the proof of similar laws to multi-model ensemble scenarios for wider applications. Taking the dual model scenario as an example is merely for the convenience of demonstrating the properties and experiments.
@@ 5 Experiments
This article chooses the splitting point based on AR law and conducts experimental verification using the introduced simple ensemble method. In the field of NLP, we designed experiments considering existing mainstream scaling-law-based LMs and custom-built LMs based on prevalent architectures. The aim is to verify the universality of our approach. First, we tested our method on several existing scaling law-based model series (e.g., GPT-2 33 ). Second, we performed ablation studies using a model built on the Transformer 31 architecture, deliberately avoiding task-specific design modifications and comparing with existing models. Finally, we retrained models of varying scales based on the mainstream architectures LSTM 29 ; 30 , Transformer, and GRU 32 for further experimentation. This chapter demonstrates that the proposed framework is universally applicable to various models, including mainstream models, custom models, and combinations thereof. Furthermore, based on the principle of relative overfitting, we validated the AR framework in computer vision (CV) and AI for science.
@@ 5.1 Mainstream Language Model Testing
This paper conducted experiments using the proposed AR framework on several pre-trained models. The tests were performed using GPT-2 33 on fundamental language modeling benchmarks including WikiText2, WikiText103 34 , text8 14 , Pile 39 ; 40 (using a randomly sampled subset of the Pile test set), and 1BW 38 , as well as on the long-context benchmark LAMBADA 35 and the QA benchmarks CBT 36 and the subject examination benchmark ARC 37 . We also evaluated the performance of models with different architectures: Pythia 41 (multiscale scaling law), RWKV 42 (combining RNN and Transformer), Mamba 43 (state-space model series), and Qwen 44 ; 45 (commercial LLM) - using the AR framework for basic language modeling tasks. We used the LAMBADA dataset for the language modeling evaluation (perplexity) and its original long-context dependency evaluation to compare the behavior under the AR framework between general language modeling and specific tasks.
The results in Table 1 demonstrate that all models achieve universal performance improvements through our AR framework. We can achieve superior performance in many scenarios with significantly fewer parameters and less computation time. For instance, on LAMBADA (PPL), Pile, and other benchmarks, our method consistently outperforms the strategy of simply increasing the LLM’s parameter count, achieving superior parameters, computation time, and performance. This advantage becomes more pronounced as the model parameters increase. For example, based on the improvement rate observed when scaling GPT-2 from 762M to 1.5B, and roughly estimating the improvement from 1.5B to 3B using the power law model proposed in scaling laws, the performance of the 1.5B+762M combination almost universally surpasses the projected 3B model across all benchmarks. Notably, our method’s advantages become more significant with larger models, while the linear complexity related to vocabulary size becomes negligible compared to increasing the parameter count. Furthermore, as the main model’s parameters grow, the magnitude of performance improvement increases rather than diminishes, making us optimistic about its performance on even larger models and hopeful that it can contribute to alleviating the scaling law bottleneck.
Table 1: The table shows the test results of the classic model GPT2 based on the AR framework with four parameter sizes from 117M to 1.5B. LTD is the long context benchmark under the LAMBADA dataset. The results of RWKV, Mamba, Pythia, and Qwen can be found in the Appendix D.1.
@@ 5.2 Ablation Experiments with Custom-Built Models Based on Mainstream Architectures
In this section, we use the vocabularies corresponding to Pythia, RWKV, Mamba, and GPT-2 to verify the universality and flexibility of our proposed method. Without incorporating any specific design elements, we employed a simple Transformer architecture to train a reference model with only approximately 180M parameters. In practice, the standalone performance of this model is not particularly strong.
We better demonstrate the regularities of ensemble theory under the AR framework through the design of experiments. In WikiText103 and LAMBADA datasets, we performed predictions using only the model trained on WikiText2, this is to illustrate that fine-grained segmentation under the AR framework allows even poorly performing models to have a positive influence on better models rather than a negative one. This was done to confirm that the performance improvement achieved by the AR framework is not merely due to having "seen the data". Furthermore, to verify the effectiveness of our method, we trained models on the 1BW and text8 datasets using their respective training sets to achieve substantial performance gains. Training different models (with poor, moderate, and good performance) based on different datasets aims to better observe the relative overfitting and performance of the AR framework under various circumstances. In addition, we will present the results based on the same model in Appendix D.2.
The results in Table 2 indicate that, through appropriate segmentation, even subpar models can still exert a positive influence. Average models are capable of significantly enhancing the performance of the primary model, while the positive impact generated by superior models is quite remarkable.
@@
Figure 2: Test results of AR framework based on simple LSTM, Transformer and GRU architectures under different embedding dimensions of SLM and LLM.
@@ 5.3 Self-built model testing based on mainstream architectures
This paper conducted tests based on the current mainstream model architectures LSTM 29 ; 30 , Transformer 31 , and GRU 32 , utilizing the same vocabulary but varying embedding dimensions. In practice, we used very simple architectures 49 ; 50 ; 51 ; 52 without any specific modifications to the model design. The ensemble theory we proposed is still valid (as illustrated in Figure 2).
@@ 5.4 Experiments in the rest of the machine learning field
Based on the principles of relative overfitting and the AR framework, we employed a similar methodology to conduct experiments with our approach across a broader range of machine learning domains, including CV and AI for Science, thus validating the general effectiveness of our proposed framework. We conducted experiments using different sizes of ResNet64 models and ESM65 models on image classification and protein sequence prediction tasks (in the Appendix C).
5.5 AR law
In order to illustrate the universality of the ensemble theory we have proposed, this section presents a portion of the results from over ten thousand experiments (as illustrated in Figure 3), while another portion is allocated for more detailed weight searches. The remaining results of the organized section can be seen in the appendix.
Figure 3: AR law in the GPT2 model combination.
@@ 6 Discussion and Future work
  • Model Selection Trade-offs: For homogeneous scaling models, while performance showed stability (no significant fluctuations) as the main model grew, the magnitude of improvement decreased as the parameter ratio expanded. Larger ratios required more extreme leverage, reducing the effectiveness of SLM and potentially limiting resource savings. In specific domains, nonhomogeneous models are a better choice.
  • Better Improvement Magnitude: Although this paper proposes the AR law based on relative overfitting and demonstrates its universal, stable, and practical improvements along with good flexibility under a suitable threshold point, significant room for progress remains beyond the decent performance improvements already achieved. To reveal fundamental principles, our method achieved these results despite lacking any specific model design, highlighting the potential yet to be unlocked. Therefore, with a deeper understanding of the entire system and relative overfitting, key research questions worth exploring include different combination methods based on the AR law, the selection and innovation of low-parameter reference model architectures serving as backbones, and how to maintain the information purity of the SLM better. We anticipate that addressing these questions will lead to greater magnitudes of improvement.
@@
  • Broader Applicability: Although this paper focuses mainly on the NLP domain and has also conducted some experiments in CV and AI for science, a broader discussion in more fields was not feasible due to limited time and resources. The applicability of the relative overfitting phenomenon and the proposed AR framework in a broader range of domains warrants further exploration.